two trains left a station at the same time one traveled north at twice th speed after 3 hours the trains were 400 miles apart how fast was each train traveling
Answer by mananth(16946)

one train speed = x
other train 2x
add up the speeds
=3x
distance =400
time = 3
d/r=t
400/3x= 3
9x=400
/9
x=44.44 mph
other train 88.88 mph